OpenStudy (anonymous):

If csc x= 4(square root of 3)/3 and cos x<0, sin x is what: A) -(sqrt13)/4, B)-(4sqrt13)/13, C) sqrt3/4, D)-sqrt39/13, E)-sqrt39/3

OpenStudy (anonymous):

csc x = 1/sin x => sin x = 1/csc x = 3/4*sqrt3 = 3sqrt3/12 cos x = sqrt(1 - 27/144) cos x = -3sqrt13/12

OpenStudy (anonymous):

im sorry, can u simplify that for me i dont understand the answer thnx